40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3585     Entry time: Fri Sep 17 17:35:35 2010
Author: steve 
Type: Update 
Category: PEM 
Subject: cleaned up at 1Y1 PSL rack 

Cleaned up cables on the top and bottom. Vacuumed both areas. We still have some remaining shading from the MOPA umbilical and more unknown BNC cables hanging around.

ELOG V3.1.3-